OpenStudy (anonymous):

find the value of theta for 0

OpenStudy (phi):

no can you solve for x in \[ x^3 = -\frac{1}{8} \]?

OpenStudy (anonymous):

2^-3

OpenStudy (phi):

you need to review cube roots x= -½ (because -½ * -½ * -½ = -⅛ that means in your problem sin theta= -½ now find the "inverse sin" but you (should) know sin 30 deg = ½

OpenStudy (phi):

in case you did not know \[ \sin^3 \theta \text{ means } ( \sin \theta)^3 \]

OpenStudy (phi):

I always associate "sin" with "y" in the unit circle -½ means the y value is negative. there are two quadrants with a negative y